# Leveraging Nested MLMC for Sequential Neural Posterior Estimation with Intractable Likelihoods

### v1

@@ 3.1 Basic idea of MLMC
Nested simulation combined with the MLMC method has been studied for other applications [18, 19, 22, 30]. In this paper, MLMC is used to derive an unbiased estimator for APT. Increasing $M$ for $\rho_{\phi,M}$ eventually recovers the desired $\rho_{\phi}$, this allows the exchange of expectation and limitation that
$\lim_{M\to\infty}\mathbb{E}\left[\rho_{\phi,M}\right]=\mathbb{E}\left[\rho_{\phi}\right]=\nabla_{\phi}\mathcal{L}(\phi).$ (16)
Then we take a monotonically increasing sequence for inner sample size $\{M_{\ell}\}_{\ell=1}^{\infty}$ with $M_{\ell}=2^{\ell}M_{0}=2M_{\ell-1}$, then (16) can be reformulated as
$\nabla_{\phi}\mathcal{L}(\phi)=\mathbb{E}\left[\rho_{\phi,M_{0}}\right]+\sum_{\ell=1}^{\infty}\mathbb{E}\left[\rho_{\phi,M_{\ell}}-\rho_{\phi,M_{\ell-1}}\right].$ (17)
Set $\{\Delta\rho_{\phi,\ell}\}_{\ell=0}^{\infty}$ to be a sequence of random variables satisfying
$\mathbb{E}\left[\Delta\rho_{\phi,0}\right]=\mathbb{E}\left[\rho_{\phi,M_{0}}\right],\quad\mathbb{E}\left[\Delta\rho_{\phi,\ell}\right]=\mathbb{E}\left[\rho_{\phi,M_{\ell}}-\rho_{\phi,M_{\ell-1}}\right]\quad(\ell\geq 1),$
then (17) can be reformulated in terms of the infinite sum of the expected values of $\{\Delta\rho_{\phi,\ell}\}_{\ell=0}^{\infty}$, i.e., $\nabla_{\phi}\mathcal{L}(\phi)=\sum_{\ell=0}^{\infty}\mathbb{E}\left[\Delta\rho_{\phi,\ell}\right]$. In the case of the loss function, we also introduce the similar sequence $\{\Delta\psi_{\phi,\ell}\}_{\ell=0}^{\infty}$ satisfying
$\mathbb{E}\left[\Delta\psi_{\phi,0}\right]=\mathbb{E}\left[\psi_{\phi,M_{0}}\right],\quad\mathbb{E}\left[\Delta\psi_{\phi,\ell}\right]=\mathbb{E}\left[\psi_{\phi,M_{\ell}}-\psi_{\phi,M_{\ell-1}}\right]\quad(\ell\geq 1),$ (18)
similarly we also have $\mathcal{L}(\phi)=\sum_{\ell=0}^{\infty}\mathbb{E}\left[\Delta\psi_{\phi,\ell}\right]$.
3.1.1 Antithetic construction
Motivated by the prevailing literature on MLMC [20, 8, 19, 22], this paper presents the construction of the sequence $\{\Delta\rho_{\phi,\ell}\}_{\ell=0}^{\infty}$ with antithetic coupling, leading to a faster rate of convergence for smooth functions. The fundamental concept of antithetic coupling involves the selection of two non-overlapping subsets, each with a size of $M_{\ell-1}$ from the $M_{\ell}$ inner samples $\{\theta_{j}^{\prime}\}_{j=1}^{M_{\ell}}$ utilized for computing $\rho_{\phi,M_{\ell}}$ given in (13). This results in two independent realizations of $\rho_{\phi,M_{\ell-1}}$, identified as $\rho^{(a)}_{\phi,M_{\ell-1}}$ and $\rho^{(b)}_{\phi,M_{\ell-1}}$. Specifically,
@@ 3.1.1 Antithetic construction
The notations $\rho^{(b)}_{\phi,M_{\ell-1}}$ , $\nabla g^{(b)}_{\phi,M_{\ell-1}}(x)$, and $g^{(b)}_{\phi,M_{\ell-1}}(x)$ are defined in the similar way by using $\theta^{\prime}_{j}$, $j=M_{\ell-1}+1,\dots,M_{\ell}$ instead. Then the antithetic construction of $\{\Delta\rho_{\phi,\ell}\}_{\ell=0}^{\infty}$ is then given as
@@
The antithetic construction of $\{\Delta\psi_{\phi,\ell}\}_{\ell=0}^{\infty}$ can be similarly given.
3.1.2 Variance of multilevel Monte Carlo estimator
In this section, we will give a series of detailed evaluation of the variance of $\Delta\rho_{\phi,\ell},\rho_{\phi,M_{\ell}}$ and $\Delta\psi_{\phi,\ell}$ which will serve as important tools when we evaluate the variance of some other MLMC methods in the coming sections. Before we present the main results, we firstly introduce the following key lemma which is stated as Lemma 1 in [19].
Lemma 3.1.
Let $X$ be a real-valued random variable with mean zero, and let $\overline{X}_{N}$ be an average of $N$ i.i.d samples of $X$. If for $\mathbb{E}[|X|^{u}]<\infty$ for $u\geq 2$, there exists a constant $C_{u}>0$ depending only on $u$ such that
for any $c>0$.
The following theorem studies the expectation and variance of $\Delta\psi_{\phi,\ell}$ as $\ell\to\infty$. In this paper, the notation $\lesssim$ is used to signifies the omission of a constant that is irrelevant to the asymptotic behavior.
@@ 3.1.2 Variance of multilevel Monte Carlo estimator
If there exist $s,z>2$ with $(s-2)(z-2)\geq 4$ such that for any $\phi$,
$\quad\mathbb{E}_{x,\theta^{\prime}}\left[\left|\frac{g_{\phi}(x,\theta^{\prime})}{Z(x,\phi)}\right|^{s}\right]<\infty\quad and\quad\mathbb{E}_{x,\theta^{\prime}}\left[\left|\log\frac{g_{\phi}(x,\theta^{\prime})}{Z(x,\phi)}\right|^{z}\right]<\infty,$
we have
$\mathrm{E}\left[\|\Delta\psi_{\phi,\ell}\|_{2}^{2}\right]\lesssim M_{\ell}^{-r_{1}},$
where $M_{\ell}=M_{0}2^{\ell}$, $r_{1}=\min(s(z-2)/2z,2)$.
@@
This proof follows an argument similar to Theorem 2 in [22]. ∎
When stochastic gradient descent method is applied, it is also desirable to examine the expectation and variances of the $\ell_{2}$ norm of the gradient. To this end, we establish the following theorem based on the work of [23]. The proof of this theorem is detailed in Appendix A.1.
@@
Assume that there exists $s>2$ such that for any $\phi$
$\underset{x,\phi,\theta}{\sup}\|\nabla_{\phi}\log g_{\phi}(x,\theta)\|_{\infty}<\infty,\,\mathbb{E}_{x,\theta}\left[\left|\frac{g_{\phi}(x,\theta)}{Z(x,\phi)}\right|^{s}\right]<\infty,$
we then have
$\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|_{2}\right]\lesssim M_{\ell}^{-1},\quad\mathrm{E}\left[\|\Delta\rho_{\phi,\ell}\|_{2}^{2}\right]\lesssim M_{\ell}^{-r_{2}},$
where $M_{\ell}=M_{0}2^{\ell}$, $r_{2}=\min(s,4)/2\in(1,2]$.
For the convergence of $\rho_{\phi,M_{\ell}}$, we establish the following theorem.
Theorem 3.4.
Under the setting of Theorem 3.3, we have $\mathrm{Var}\left[\rho_{\phi,M_{\ell}}\right]\lesssim M_{\ell}^{-1}.$
A detailed version and proof of this theorem is provided in Appendix A.2. This theorem can be similarly extended to the case where arbitrarily number of inner samples are used for the estimation of the gradient. As a result, it indicates the convergence rate of the gradient estimator is the same as the loss function.
@@ 3.2.1 RU-MLMC
Recall that both loss function and its gradient can be represented as the summation of the expected value of the random variable sequence, and this topic has been well studied. We introduce a non-negative integer-valued random variable $L$ that independent of the two random variable sequences, with its probability mass function $\mathbb{P}(L=\ell)=w_{\ell}$. We then have
@@
leads to an unbiased Monte Carlo estimator of the gradient $\nabla_{\phi}\mathcal{L}(\phi)$. We similarly define $U_{\mathrm{RU}}={\omega_{L}}^{-1}{\Delta\psi_{\phi,L}}$ for the query of the estimation of the loss function $\mathcal{L}(\phi)$. This method is known as randomized unbiased multilevel Monte Carlo (RU-MLMC) method [51]. In this study, we take $L$ as a geometric distribution $\mathrm{Ge}(p)$ with $\omega_{\ell}=(1-p)^{\ell}p$ and $p=1-2^{-\alpha}$. To ensure a finite variance and finite expected computational cost for $V_{\mathrm{RU}},U_{\mathrm{RU}}$, it is required that $\alpha\in(1,\min\left(r_{1},r_{2}\right))$. This is further examined with the following theorem.
@@
Under the setting of Theorems 3.2 and 3.3, let $L\sim\mathrm{Ge}(p)$ with $p=1-2^{-\alpha}$ and $1<\alpha<\min\left(r_{1},r_{2}\right)$, where $r_{1}$ and $r_{2}$ are from Theorems 3.2 and 3.3 respectively. Then we have
@@
where the constants $A,B$ are independent of $\alpha$ and $r_{1}$.
@@
The proof of this theorem follows in line of Theorem 3.6. ∎
This theorem indicates that a large $\alpha$ leads to less expected total computational burden, but larger variance for the estimators of loss and its gradient.

### v2

@@ 3.1 Basic idea of MLMC
We first introduce the idea of MLMC in a generic setting following [16]. Consider the problem of estimating $\mathbb{E}[P]$, where the sampling of $P$ is costly or even infeasible. Suppose that $P$ can be approximated by tractable $P_{\ell}$ with increasing accuracy, but also increasing cost, as $\ell\to\infty$. As a result, the expectation $\mathbb{E}[P]$ can be estimated by the simple Monte Carlo (i.e., single level) estimator
$\hat{\mu}_{L,N}:=\frac{1}{N}\sum_{i=1}^{N}P_{L}^{(i)},$ (17)
where $P_{L}^{(i)}$ are iid copies of $P_{L}$. Assume that $P_{\ell}$ has a bias $|\mathbb{E}[P_{\ell}-P]|=\mathcal{O}(2^{-\alpha\ell})$ and an expected cost $\mathcal{O}(2^{\gamma\ell})$, where $\alpha,\gamma>0$. The estimator $\hat{\mu}_{L,N}$ then has an MSE of $\mathcal{O}(2^{-2\alpha L})+\mathcal{O}(1/N)$. To ensure an MSE of $\mathcal{O}(\epsilon^{2})$, one can take $L=\mathcal{O}(\log(\epsilon^{-1/\alpha}))$ and $N=\mathcal{O}(\epsilon^{-2})$, yielding an average computational cost of $\mathcal{O}(N2^{\gamma L})=\mathcal{O}(\epsilon^{-2-\gamma/\alpha})$. MLMC can greatly reduce the computational cost by performing most simulations with low accuracy (low level) at a low cost, with relatively few simulations being performed at high accuracy (high level) and a high cost. Specifically, using the simple identity
$\mathbb{E}[P_{L}]=\mathbb{E}[P_{0}]+\sum_{\ell=1}^{L}\mathbb{E}[P_{\ell}-P_{\ell-1}]=\sum_{\ell=0}^{L}\mathbb{E}[\Delta P_{\ell}],$
where $\mathbb{E}[\Delta P_{\ell}]=\mathbb{E}[P_{\ell}-P_{\ell-1}]$ for $\ell\geq 1$ and $\mathbb{E}[\Delta P_{0}]=\mathbb{E}[P_{0}]$, we can use the following unbiased estimator for $\mathbb{E}[P_{L}]$,
$\hat{\mu}_{\mathrm{mlmc}}:=\sum_{\ell=0}^{L}\left(\frac{1}{N_{\ell}}\sum_{i=1}^{N_{\ell}}\Delta P_{\ell}^{(i)}\right),$ (18)
where $\Delta P_{\ell}^{(i)}$ are iid copies of $\Delta P_{\ell}$, the sample sizes $N_{\ell}$ decrease with level $\ell$.
Now suppose that $\Delta P_{\ell}$ has a variance of $\mathcal{O}(2^{-r\ell})$ with an average cost of $\mathcal{O}(2^{\gamma\ell})$. For a given MSE of $\mathcal{O}(\epsilon^{2})$, [16] showed that choosing proper finest level $L$ and sample sizes $N_{\ell}$ render a computational complexity $C$ with bound
$\mathbb{E}[C]=\begin{cases}\mathcal{O}(\epsilon^{-2})&r>\gamma,\\ \mathcal{O}(\epsilon^{-2}(\log\epsilon)^{2})&r=\gamma,\\ \mathcal{O}(\epsilon^{-2-(\gamma-\beta)/\alpha})&r<\gamma.\end{cases}$ (19)
As stated before, it requires a computational complexity of $\mathcal{O}(\epsilon^{-2-\gamma/\alpha})$ for the simple Monte Carlo to achieve the same level of MSE. From the view point of complexity, MLMC beats the simple Monte Carlo (i.e., the single level Monte Carlo) for all cases of $r$ and achieves the optimal complexity of $\mathcal{O}(\epsilon^{-2})$ when $r>\gamma$.
Nested simulation combined with the MLMC method has been studied for other applications [17, 18, 21, 29]. We next show how to use MLMC to estimate the gradient $\nabla_{\phi}\mathcal{L}(\phi)=\mathbb{E}\left[\rho_{\phi}\right]$, which is a nested expectation. To fit the MLMC setting, we can take $P=\rho_{\phi}$ and $P_{\ell}=\rho_{\phi,M_{\ell}}$, where $M_{\ell}=2^{\ell}M_{0}=2M_{\ell-1}$. It is clear that $P_{\ell}$ converges $P$ as $\ell\to\infty$. The core of MLMC is to construct coupling estimators $\Delta P_{\ell}$ with a variance decaying as $\mathcal{O}(2^{-r\ell})$. To gain a large $r$, we use the antithetic coupling construction as in [8, 18, 19, 21], which is better than the naive coupling $\Delta P_{\ell}=\rho_{\phi,M_{\ell}}-\rho_{\phi,M_{\ell-1}}$. The fundamental concept of antithetic coupling involves the selection of two non-overlapping subsets, each with a size of $M_{\ell-1}$ from the $M_{\ell}$ inner samples $\{\theta_{j}^{\prime}\}_{j=1}^{M_{\ell}}$ utilized for computing $\rho_{\phi,M_{\ell}}$ given in (14). This results in two independent realizations of $\rho_{\phi,M_{\ell-1}}$, identified as $\rho^{(a)}_{\phi,M_{\ell-1}}$ and $\rho^{(b)}_{\phi,M_{\ell-1}}$. Specifically,
@@
The notations $\rho^{(b)}_{\phi,M_{\ell-1}}$, $\nabla g^{(b)}_{\phi,M_{\ell-1}}(x)$, and $g^{(b)}_{\phi,M_{\ell-1}}(x)$ are defined in the similar way by using $\theta^{\prime}_{j}$, $j=M_{\ell-1}+1,\dots,M_{\ell}$ instead. Define
@@
It suffices to take $\Delta P_{\ell}=\Delta\rho_{\phi,\ell}$ since
$\mathbb{E}\left[\Delta\rho_{\phi,0}\right]=\mathbb{E}\left[\rho_{\phi,M_{0}}\right],\quad\mathbb{E}\left[\Delta\rho_{\phi,\ell}\right]=\mathbb{E}\left[\rho_{\phi,M_{\ell}}-\rho_{\phi,M_{\ell-1}}\right]\quad(\ell\geq 1).$
In the case of estimating the loss function $\mathcal{L}(\phi)=\mathbb{E}[\psi_{\phi}]$, the antithetic construction of the coupling estimators $\{\Delta\psi_{\phi,\ell}\}_{\ell=0}^{\infty}$ can be similarly given by
$\Delta\psi_{\phi,0}=\psi_{\phi,M_{0}},\quad\Delta\psi_{\phi,\ell}=\psi_{\phi,M_{\ell}}-\frac{1}{2}\left(\psi^{(a)}_{\phi,M_{\ell-1}}+\psi^{(b)}_{\phi,M_{\ell-1}}\right)\quad(\ell\geq 1),$ (21)
where $\psi_{\phi,M_{\ell}}$ is defined in (13), and
$\psi_{\phi,M_{\ell-1}}^{(a)}=\log\frac{1}{M_{\ell-1}}\sum_{j=1}^{M_{\ell-1}}\frac{g_{\phi}(x,\theta^{\prime}_{j})}{g_{\phi}(x,\theta)},\quad\psi_{\phi,M_{\ell-1}}^{(b)}=\log\frac{1}{M_{\ell-1}}\sum_{j=M_{\ell-1}+1}^{M_{\ell}}\frac{g_{\phi}(x,\theta^{\prime}_{j})}{g_{\phi}(x,\theta)}.$
We have $\gamma=1$ for $\Delta\rho_{\phi,\ell}$ and $\Delta\psi_{\phi,\ell}$. We next examine the values of $\alpha$ and $r$. To this end, the next two theorems provide upper bounds for the first and second moments of $|\Delta\psi_{\phi,\ell}|$ and $\left\|\Delta\rho_{\phi,\ell}\right\|_{2}$ as $\ell\to\infty$.
@@
If there exist $p,q>2$ with $(p-2)(q-2)\geq 4$ such that for any $\phi\in\Phi$,
$\quad\mathbb{E}\left[\left(\frac{g_{\phi}(x,\theta)}{Z(x,\phi)}\right)^{p}\right]<\infty\quad and\quad\mathbb{E}\left[\left|\log\frac{g_{\phi}(x,\theta)}{Z(x,\phi)}\right|^{q}\right]<\infty,$
where $(\theta,x)\sim\tilde{p}(\theta)p(x|\theta)$, we have
$\mathbb{E}\left[|\Delta\psi_{\phi,\ell}|\right]=\mathcal{O}(2^{-\ell}),\quad\mathbb{E}\left[\Delta\psi_{\phi,\ell}^{2}\right]=\mathcal{O}(2^{-r_{1}\ell}),$
where $r_{1}=\min(p(q-2)/(2q),2)\in(1,2]$.
@@
This proof follows an argument similar to Theorem 2 in [21], which particularly showed that $\mathbb{E}\left[|\Delta\psi_{\phi,\ell}|\right]=\mathcal{O}(2^{-\alpha_{1}\ell})$ with $\alpha_{1}=\min(p(q-1)/(2q),1)$, Since $(p-2)(q-2)\geq 4$, then we actually have $\alpha_{1}=1$. ∎
@@
If there exists $s>2$ such that for any $\phi\in\Phi$,
$\underset{x,\phi,\theta}{\sup}\|\nabla_{\phi}\log g_{\phi}(x,\theta)\|_{\infty}<\infty,\quad\mathbb{E}\left[\left(\frac{g_{\phi}(x,\theta)}{Z(x,\phi)}\right)^{s}\right]<\infty,$
where the expectation is taken under the law $(\theta,x)\sim\tilde{p}(\theta)p(x|\theta)$, we then have
$\mathbb{E}\left[\left\|\Delta\rho_{\phi,\ell}\right\|_{2}\right]=\mathcal{O}(2^{-\ell}),\quad\mathbb{E}\left[\|\Delta\rho_{\phi,\ell}\|_{2}^{2}\right]=\mathcal{O}(2^{-r_{2}\ell}),$
where $r_{2}=\min(s,4)/2\in(1,2]$.
Theorem 3.2 is based on the work of [22] and its proof is detailed in Appendix A. By Theorem 3.1 and Theorem 3.2 , $\Delta\psi_{\phi,\ell}$ and $\Delta\rho_{\phi,\ell}$ have a variance of $\mathcal{O}(2^{-r\ell})$ with $r>\gamma=1$. Using the fact that
$|\mathbb{E}[P_{\ell}-P]|=\left|\sum_{\ell^{\prime}=\ell+1}^{\infty}\mathbb{E}[\Delta P_{\ell}]\right|\leq\sum_{\ell^{\prime}=\ell+1}^{\infty}\mathbb{E}[\left|\Delta P_{\ell}\right|],$
we have the bias rate $\alpha=1$ for both cases. According to (19), the resulting MLMC estimator (18) yields the optimal complexity $\mathcal{O}(\epsilon^{-2})$. We should note that the nested APT estimator has a complexity of $\mathcal{O}(\epsilon^{-2-\gamma/\alpha})=\mathcal{O}(\epsilon^{-3})$. Although the MLMC estimator (18) gains lower complexity than the nested APT estimator, it is biased and depends on the prespecified error threshold $\epsilon$. Moreover, the sample sizes $N_{\ell}$ for each level and the finest level $L$ need to be tuned by the threshold $\epsilon$. For practical training problems, it may be difficult to choose the threshold $\epsilon$. Since MLMC achieves the optimal regime $r>\gamma$, one can design unbiased MLMC estimators for eliminating the bias of the MLMC estimator (18). The unbiased MLMC also admits the optimal complexity $\mathcal{O}(\epsilon^{-2})$ without choosing the sample sizes $N_{\ell}$.
@@ 3.2 Unbiased MLMC for APT
In this subsection, we develop a series of unbiased estimators for the loss function $\mathcal{L}(\phi)$ and its gradient $\nabla_{\phi}\mathcal{L}(\phi)$ based on the MLMC approach.
@@ 3.2.1 RU-MLMC
Recall that both the loss function and its gradient can be represented as the summation of the expected value of the random variable sequence, and this topic has been well studied. We introduce a non-negative integer-valued random variable $L$ that is independent of the two random variable sequences, with its probability mass function $\mathbb{P}(L=\ell)=w_{\ell}$. By the law of total expectation, we have
@@
leads to an unbiased Monte Carlo estimator of the gradient $\nabla_{\phi}\mathcal{L}(\phi)$.
We similarly define
$U_{\mathrm{RU}}={\omega_{L}}^{-1}{\Delta\psi_{\phi,L}}$
for the query of the estimation of the loss function $\mathcal{L}(\phi)$. This method is known as the randomized unbiased multilevel Monte Carlo (RU-MLMC) method [51]. In this study, we take $L$ as a geometric distribution $\mathrm{Ge}(p)$ with $\omega_{\ell}=(1-p)^{\ell}p$ and $p=1-2^{-\alpha}$. To ensure a finite variance and finite expected computational cost for $V_{\mathrm{RU}},U_{\mathrm{RU}}$, it is required that $\alpha\in(1,\min\left(r_{1},r_{2}\right))$, where $r_{1}$ and $r_{2}$ are from Theorem 3.1 and Theorem 3.2, respectively. This is further examined with the following theorem. Since the gradient $V_{\mathrm{RU}}$ is a vector, the notation $\mathrm{Var}[V_{\mathrm{RU}}]$ used in the following applies to the variance of each component of $V_{\mathrm{RU}}$ for simplicity. Denote the cost of RU-MLMC as $\mathrm{Cost}_{\mathrm{RU}}$.
@@
Consider the settings of Theorem 3.1 and Theorem 3.2. If $L\sim\mathrm{Ge}(p)$ with $p=1-2^{-\alpha}$ and $1<\alpha<\min\left(r_{1},r_{2}\right)$, then we have
@@
where the constants $A,B$ are independent of $\alpha$, $r_{1}$ and $r_{2}$.
@@
It is sufficient to bound the second moments of $U_{\mathrm{RU}}$ and $V_{\mathrm{RU}}$. First, by taking $w_{\ell}=(1-p)^{\ell}p=(1-2^{-\alpha})2^{-\alpha\ell}$ and using Theorem 3.1, we have
$\mathrm{Var}\left[U_{\mathrm{RU}}\right]$ $\leq\mathbb{E}\left[\left(\frac{\Delta\psi_{\phi,L}}{w_{L}}\right)^{2}\right]=\sum_{\ell=0}^{\infty}\frac{\mathbb{E}\left[\Delta\psi_{\phi,\ell}^{2}\right]}{w_{\ell}}\leq\frac{A}{(1-2^{-\alpha})}\sum_{\ell=0}^{\infty}2^{(\alpha-r_{1})\ell}$
$=\frac{A}{(1-2^{\alpha-r_{1}})(1-2^{-\alpha})}.$
Bounding $\mathrm{Var}[V_{\mathrm{RU}}]$ is just a replica by using Theorem 3.2.
As for the average cost, we have the following
$\mathrm{Cost}_{\mathrm{RU}}\propto\sum_{\ell=0}^{\infty}w_{\ell}M_{\ell}=M_{0}(1-2^{-\alpha})\sum_{\ell=0}^{\infty}2^{(1-\alpha)\ell}=M_{0}\frac{2^{\alpha}-1}{2^{\alpha}-2}.$
Theorem 3.3 indicates that a large $\alpha$ leads to less expected total computational burden, but larger variance for the estimators of loss and its gradient.